# Robust Kernel Hypothesis Testing under Data Corruption

### v1

@@ Abstract
We propose two general methods for constructing robust permutation tests under data corruption. The proposed tests effectively control the non-asymptotic type I error under data corruption, and we prove their consistency in power under minimal conditions. This contributes to the practical deployment of hypothesis tests for real-world applications with potential adversarial attacks. One of our methods inherently ensures differential privacy, further broadening its applicability to private data analysis. For the two-sample and independence settings, we show that our kernel robust tests are minimax optimal, in the sense that they are guaranteed to be non-asymptotically powerful against alternatives uniformly separated from the null in the kernel MMD and HSIC metrics at some optimal rate (tight with matching lower bound). Finally, we provide publicly available implementations and empirically illustrate the practicality of our proposed tests.
Robust Kernel Hypothesis Testing
under Data Corruption
Antonin Schrab
Centre for Artificial Intelligence
Gatsby Computational Neuroscience Unit
University College London & Inria London
a.schrab@ucl.ac.uk
Ilmun Kim
Department of Statistics & Data Science
Department of Applied Statistics
Yonsei University
ilmun@yonsei.ac.kr
May 30, 2024
$\star$\$$\star$\$footnotetext: Both authors contributed equally to this work.
Abstract
@@ 1 Introduction
Robust testing. There is an extensive body of work on robust testing developed under various contamination models. One popular framework is Huber’s $\epsilon$-contamination model [35] where the observed data are assumed to be drawn from a mixture distribution $(1-\epsilon)P+\epsilon G$ with $P$ as the target distribution and $G$ as the contamination distribution. Our work is concerned with a stronger adversarial contamination model where an adversary selects $r$ observations out of $n$ data points and replace them with arbitrary values. Under different contamination models, there has been a flurry of recent work on robust testing such as the two-point testing problem [15, 42], covariance testing [19], mean testing [20, 27, 13], identity testing [3]. See [18] for a recent survey on robust statistics. We aim to advance the field by studying robust testing for exchangeability (see Section 2.1 for a formal setup), and establishing minimax testing rates in terms of the kernel metrics under data corruption.
Differential privacy. Differential privacy [23] is a popular framework for private data analysis that aims to return an output insensitive to any single individual’s data point. As discussed in [22, 7], there is a close connection between robust statistics and differential privacy. Both frameworks require that algorithms remain stable under data perturbation. Recent research has delved into extending non-private tests to privacy settings and studying their asymptotic properties [25, 17, 12, 47, 46] and non-asymptotic properties [11, 6, 5, 1, 38]. Our work is particularly motivated by the work of [38] that introduces differentially private permutation tests. Building on [38], as well as the close connection between differential privacy and robustness, we propose tests which are both differentially private and robust to data corruption.
@@
Permutation tests. As mentioned earlier, our main interest is in designing robust tests for assessing the exchangeability of data under the null hypothesis. A gold standard method for testing exchangeability is the permutation test, which leverages the permutation-invariant properties of the data distribution under the null hypothesis. This ensures rigorous control of the type I error rate as demonstrated in [e.g., 48, 34]. When the data are arbitrarily corrupted, however, the exchangeability assumption no longer holds, and naively applying the permutation test using the corrupted data can potentially inflate the type I error rate. To address this miscalibration issue, it is essential to either modify the test statistic or adjust the permutation critical value under data corruption. This process inevitably compromises the power of the test, and the key challenge is then to balance a trade-off between robustness and power. We approach this problem under the minimax testing framework as in [4, 50, 3, 38] and explain the fundamental limit of testing under data corruption.
@@
Kernel MMD and HSIC. The Maximum Mean Discrepancy (MMD) [29] and the Hilbert–Schmidt Independence Criterion (HSIC) [30] are two prominent kernel-based measures for assessing homogeneity and dependence between two random quantities. Since their introduction, there have been significant developments in related topics, including time-efficient kernel tests [32, 55, 56, 54, 51, 21], adaptive kernel selections [50, 51, 49, 33, 21, 8, 14] and minimax optimality of kernel testing [4, 43, 50, 52, 53, 33]. However, prior work has predominantly focused on standard i.i.d. data without any perturbation, potentially limiting their applicability to real-world scenarios with anomalies or adversarial attacks. While a few studies, such as [10, Section 3.3] and [16, Section 3.2.3], explore the robustness of MMD statistics in terms of estimation, there has been a limited understanding of their performance in robust testing. To fill this gap, we consider the two-sample and independence testing problems as running examples of our general methods, and propose robust tests based on the MMD and HSIC statistics. We then show that the proposed tests guarantee rigorous control of testing errors even in the presence of data corruption, and achieve minimax optimality across all levels of data corruption.
@@
Outline and summary. The remainder of this paper is organised as follows. In Section 2, we formally introduce our testing under data corruption framework, and propose two methods (Algorithm 1 and Algorithm 2) for constructing robust permutation tests from any statistic with finite sensitivity. We prove that these methods are valid (non-asymptotic control of the type I error even under data corruption) and present conditions for consistency (asymptotic control of the type II error). In Sections 3 and 4, we construct robust kernel tests for the two-sample and independence testing frameworks, respectively. For each robust test, we prove minimax optimality under data corruption in terms of the respective MMD and HSIC kernel metrics. We verify the validity of our robust tests in Section 5, and highlight their test power on corrupted synthetic and real-world data. Section 6 closes the paper with further discussions. To conserve space, we defer the technical proofs of the main results to the appendix.
@@
Notation. Consider two ordered sets $\mathcal{X}_{n}\coloneqq(X_{1},\ldots,X_{n})$ and $\tilde{\mathcal{X}}_{n}\coloneqq(\tilde{X}_{1},\ldots,\tilde{X}_{n})$. The Hamming distance $d_{\mathrm{ham}}(\mathcal{X}_{n},\tilde{\mathcal{X}}_{n}):=\sum_{i=1}^{n}\mathds{1}(X_{i}\neq\tilde{X}_{i})$ is the number of indices for which the entries of the two ordered sets are different. We let $[n]$ denote the set $\{1,\dots,n\}$ and $[n]_{0}\coloneqq\{0,\dots,n\}$. We define $\bm{\Pi}_{n}$ to be the set of all permutations of $[n]$. Given a permutation $\bm{\pi}\in\bm{\Pi}_{n}$, we denote by $\mathcal{X}_{n}^{\bm{\pi}}$ the permuted ordered set $(X_{\bm{\pi}(1)},\ldots,X_{\bm{\pi}(n)})$. For two sequences of real numbers $a_{n}$ and $b_{n}$, we write $a_{n}\lesssim b_{n}$ if $a_{n}\leq Cb_{n}$ for some constant $C>0$, and write $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$. The global sensitivity $\Delta_{T}$ of a statistic $T$ is defined as the maximum difference in the statistic’s output when evaluated on permuted datasets whose entries differ by at most one, that is
@@
$\Delta_{T}\coloneqq\sup_{\bm{\pi}\in\bm{\Pi}_{n}}\sup_{\begin{subarray}{c}\mathcal{X}_{n},\tilde{\mathcal{X}}_{n}\,:\,d_{\mathrm{ham}}(\mathcal{X}_{n},\tilde{\mathcal{X}}_{n})\leq 1\end{subarray}}\bigl{|}T(\mathcal{X}_{n}^{\bm{\pi}})-T(\tilde{\mathcal{X}}_{n}^{\bm{\pi}})\bigr{|}.$ (1)
@@
Given values $M_{0},\ldots,M_{B}$, the $(1-\alpha)$-quantile of the $M_{i}$’s is defined as $q_{1-\alpha}=\inf\bigl{\{}t\in\mathbb{R}:\frac{1}{B+1}\sum_{i=0}^{B}\mathds{1}(M_{i}\leq t)\geq 1-\alpha\bigr{\}}$. We let $\mathsf{Laplace}(0,1)$ denote a Laplace distribution with location and scale parameters $(0,1)$. We also let $\mathsf{Gaussian}(\mu,\sigma,d)$ denote a $d$-dimensional distribution with each dimension being independent and following a Gaussian distribution with mean $\mu$ and standard deviation $\sigma$. Similarly, $\mathsf{Geometric}(p,d)$ denotes a $d$-dimensional distribution with each dimension being independent and following a geometric distribution taking values in $\{0,1,2,\ldots\}$ with parameter $p\in(0,1)$.
@@ 2 Construction of hypothesis tests robust to data corruption
In Section 2.1, we formalise our novel framework for robust testing under data corruption. In Section 2.2, we construct robust tests by leveraging an existing differential privatisation (DP) method for permutation tests, and prove their validity and consistency under the data corruption framework. In Section 2.3, we introduce a novel procedure to construct robust tests under data corruption, which trades off the DP property for higher test power.
@@ 2.1 Robust testing framework under data corruption
Robustness. Hypothesis tests designed for this setting are robust to data corruption: under the null, manipulating up to $r$ samples will not make the test deviate from the null. This can also be thought of as enlarging the null hypothesis as it only needs to hold for $n-r$ samples of the data rather than for all $n$ samples. The number $r$ of sample manipulations to be robust to is specified by the user depending on the application. If $r=0$, we recover the standard testing framework. As $r$ increases, the test becomes more robust but less powerful (i.e., there is a trade-off between robustness and power). If $r=n$, the null would never be rejected.
Exchangeability. We restrict our attention to testing frameworks under which the exchangeability assumption holds [41, Chapter 15.2] in the non-corrupted setting, which means that, under the null, $\widetilde{\mathcal{X}}_{n}$ is exchangeable: for any permutation $\bm{\pi}\in\bm{\Pi}_{n}$, the joint distributions of $\widetilde{\mathcal{X}}_{n}$, and of $\widetilde{\mathcal{X}}_{n}^{\bm{\pi}}$, are the same. For example, the two-sample and independence testing frameworks, serving as primary applications of our proposal, satisfy this assumption.
Algorithm 1 Robust DP procedure
Inputs: Data $\mathcal{X}_{n}$, robustness $r$, level $\alpha$, statistic $T$, permutation number $B$, privacy $\varepsilon$.
Generate i.i.d. $\zeta_{0},\dots,\zeta_{B}\sim\mathsf{Laplace}(0,1)$.
Generate i.i.d. permutations $\bm{\pi}_{1},\dots,\bm{\pi}_{B}$ of $[n]$.
Set $\bm{\pi}_{0}=\mathrm{Id}$ and compute global sensitivity $\Delta_{T}$.
Compute $(1\!-\!\alpha e^{-r\varepsilon})$-quantile $q$ of $M_{0},\dots,M_{B}$.
Output: Reject $\mathcal{H}_{0}$ if $M_{0}>q$.
@@ 2.2 DP procedure: Robust test construction via differentially privacy
We assume the maximum number of corrupted samples $r$ to be fixed apriori. The aim is then to construct a robust test that is valid in the sense that it controls the type I error at the desired level $\alpha$ even when $r$ samples are arbitrarily corrupted. As mentioned earlier, differential privacy (DP) is closely related to statistical robustness [22, 7]. This connection naturally motivates us to leverage the recent advances in differential privatisation of hypothesis tests [45, 36, 38] for our data corruption setting. Consider any valid $\varepsilon$-DP hypothesis test with adjusted lower level $\alpha e^{-r\varepsilon}$ and pure DP parameter $\varepsilon>0$. Under the null, the DP group privacy property [24, Theorem 2.2] ensures that
$\sup_{P_{0}\in\mathcal{P}_{0}}\mathbb{P}_{P_{0}}\bigl{(}\textrm{reject }\mathcal{H}_{0}\,\,|\,\,r\textrm{ corrupted data}\bigr{)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ e^{r\varepsilon}\sup_{P_{0}\in\mathcal{P}_{0}}\mathbb{P}_{P_{0}}\bigl{(}\textrm{reject }\mathcal{H}_{0}\,\,|\,\,\textrm{uncorrupteddata}\bigr{)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \alpha.$ (2)
We choose to focus on the differential privatisation approach of [38] for permutation tests, as it benefits from stronger theoretical guarantees. In particular, their resulting DP tests are guaranteed to be valid at every sample size [38, Theorem 1], so the type I error control under data corruption of Equation 2 holds non-asymptotically. This is actually crucial as the corruption and privacy parameters $r$ and $\varepsilon$ usually depend on the sample size. Consequently, the adjusted level $\alpha e^{-r\varepsilon}$ also depends on the sample size, which can modify the asymptotic behaviour of type I error control. We simply refer to this test construction with the adjusted level as the DP procedure (see details in Algorithm 1). We can leverage the result of [38, Theorem 3] to guarantee the consistency of the DP test against data corruption (i.e., for a fixed alternative distribution, the test power converges to one asymptotically) under minimal conditions.
Lemma 1 (Validity & consistency of DP).
$\lim_{n\rightarrow\infty}\mathbb{P}_{P_{1}}\bigl{(}\,T({\widetilde{\mathcal{X}}_{n}})+2\zeta\Delta_{T}\varepsilon^{-1}>\leavevmode\nobreak\ T({\widetilde{\mathcal{X}}_{n}}^{\bm{\pi}})+2\zeta^{\prime}\Delta_{T}\varepsilon^{-1}\bigr{)}=1,$
2.3 DC procedure: Robust test construction against data corruption
We now propose an alternative procedure to construct tests which are robust to data corruption, without the need to inject random noise. First, recall that, given a statistic $T$, the classical permutation test is defined as rejecting the null when $T_{0}>q_{1-\alpha}(T_{0}\ldots,T_{B})$ where $T_{0}=T(\mathcal{X}_{n})$ and $T_{i}=T(\mathcal{X}_{n}^{\bm{\pi}_{i}})$, $i\in[B]$ for $B$ permutations $\bm{\pi}_{1},\ldots,\bm{\pi}_{B}\in\bm{\Pi}_{n}$. Such a test in the usual setting (no data corruption) controls the type I error at level $\alpha$ non-asymptotically [e.g., 34]. For our setting with at most $r$ corrupted samples, we instead define our DC test as rejecting the null when $T_{0}>q_{1-\alpha}(T_{0}\ldots,T_{B})+2r\Delta_{T}$ (see details in Algorithm 2). We prove that this results in a well-calibrated non-asymptotic test under $r$ data corruption, and provide sufficient conditions to guarantee its consistency.
@@ 2.3 DC procedure: Robust test construction against data corruption
$\lim_{n\rightarrow\infty}\mathbb{P}_{P_{1}}\bigl{(}\,T({\widetilde{\mathcal{X}}_{n}})\leavevmode\nobreak\ >\leavevmode\nobreak\ T({\widetilde{\mathcal{X}}_{n}}^{\bm{\pi}})+4r\Delta_{T}\bigr{)}=1,$
@@
We highlight that both Lemma 1 and Lemma 2 are applicable to any test statistic with finite global sensitivity $\Delta_{T}$. Moreover, their validity under data corruption holds in any testing framework for which exchangeability holds under the (uncorrupted) null hypothesis. We next turn to two-sample and independence kernel testing as specific applications of these general methods, and present more refined results.

### v3

@@ Abstract
We propose a general method for constructing robust permutation tests under data corruption. The proposed tests effectively control the non-asymptotic type I error under data corruption, and we prove their consistency in power under minimal conditions. This contributes to the practical deployment of hypothesis tests for real-world applications with potential adversarial attacks. For the two-sample and independence settings, we show that our kernel robust tests are minimax optimal, in the sense that they are guaranteed to be non-asymptotically powerful against alternatives uniformly separated from the null in the kernel MMD and HSIC metrics at some optimal rate (tight with matching lower bound). We point out that existing differentially private tests can be adapted to be robust to data corruption, and we demonstrate in experiments that our proposed tests achieve much higher power than these private tests. Finally, we provide publicly available implementations and empirically illustrate the practicality of our robust tests.
@@ 1 INTRODUCTION
Robust testing. There is an extensive body of work on robust testing developed under various contamination models. One popular framework is Huber’s $\epsilon$-contamination model  (Huber,, 1964) where the observed data are assumed to be drawn from a mixture distribution $(1-\epsilon)P+\epsilon G$ with $P$ as the target distribution and $G$ as the contamination distribution. Our work is concerned with a stronger adversarial contamination model where an adversary selects $r$ observations out of $n$ data points and replace them with arbitrary values. Under different contamination models, there has been a flurry of recent work on robust testing such as the two-point testing problem (Chen et al.,, 2016; Li et al.,, 2023), covariance testing (Diakonikolas and Kane,, 2021), mean testing (Diakonikolas et al.,, 2017; George and Canonne,, 2022; Canonne et al.,, 2023), identity testing (Acharya et al., 2021b, ). See Diakonikolas and Kane, (2019) for a recent survey on robust statistics. We aim to advance the field by studying robust testing for exchangeability (see Section 2.1 for a formal setup), and establishing minimax testing rates in terms of the kernel metrics under data corruption. A distinct line of work explores a different notion of robustness, where the goal is to determine whether a given dataset lies within an uncertainty set centered around an empirical distribution with uncertainty defined using various metrics (Gao et al.,, 2018; Wang and Xie,, 2022; Sun and Zou,, 2021, 2022, 2023). This approach contrasts with our focus on worst-case contamination as it seeks to account for distributional ambiguity rather than adversarial data corruption.
@@
Permutation tests. As mentioned earlier, our main interest is in designing robust tests for assessing the exchangeability of data under the null hypothesis. A gold standard method for testing exchangeability is the permutation test, which leverages the permutation-invariant properties of the data distribution under the null hypothesis. This ensures rigorous control of the type I error rate as demonstrated in Romano and Wolf, (e.g., 2005); Hemerik and Goeman, (e.g., 2018). When the data are arbitrarily corrupted, however, the exchangeability assumption no longer holds, and naively applying the permutation test using the corrupted data can potentially inflate the type I error rate. To address this miscalibration issue, it is essential to either modify the test statistic or adjust the permutation critical value under data corruption. This process inevitably compromises the power of the test, and the key challenge is then to balance a trade-off between robustness and power. We approach this problem under the minimax testing framework as in Albert et al., (2022); Schrab et al., (2023); Acharya et al., 2021b ; Kim and Schrab, (2023) and explain the fundamental limit of testing under data corruption.
@@
Kernel MMD and HSIC. The Maximum Mean Discrepancy (MMD, Gretton et al., 2012a, ) and the Hilbert–Schmidt Independence Criterion (HSIC, Gretton et al., 2005a, ) are two prominent kernel-based measures for assessing homogeneity and dependence between two random quantities. Since their introduction, there have been significant developments in related topics, including time-efficient kernel tests (Gretton et al., 2012b, ; Zaremba et al.,, 2013; Zhao and Meng,, 2015; Yamada et al.,, 2019; Schrab et al., 2022b, ; Domingo-Enrich et al.,, 2023), adaptive kernel selections (Schrab et al.,, 2023; Schrab et al., 2022b, ; Schrab et al., 2022a, ; Hagrass et al.,, 2024; Domingo-Enrich et al.,, 2023; Biggs et al.,, 2024; Chatterjee and Bhattacharya,, 2023) and minimax optimality of kernel testing (Albert et al.,, 2022; Li and Yuan,, 2024; Schrab et al.,, 2023; Shekhar et al.,, 2022, 2023; Hagrass et al.,, 2024). However, prior work has predominantly focused on standard i.i.d. data without any perturbation, potentially limiting their applicability to real-world scenarios with anomalies or adversarial attacks. While there exist some studies on the robustness of MMD statistics in terms of estimation (Briol et al.,, 2019, Section 3.3 and Chérief-Abdellatif and Alquier,, 2022, Section 3.2.3) and in terms of Bayesian inference (Chérief-Abdellatif and Alquier,, 2020; Dellaporta et al.,, 2022; Dellaporta and Damoulas,, 2023; Legramanti et al.,, 2025; Fazeli-Asl et al.,, 2024), there has been limited understanding of their performance in robust testing. To fill this gap, we consider the two-sample and independence testing problems as running examples of our general methods, and propose robust tests based on the MMD and HSIC statistics. We then show that the proposed tests guarantee rigorous control of testing errors even in the presence of data corruption, and achieve minimax optimality across all levels of data corruption.
@@
Outline and summary. The remainder of this paper is organised as follows. In Section 2, we formally introduce our testing under data corruption framework, and propose our DC procedure (Algorithm 1) for constructing robust permutation tests from any statistic with finite sensitivity. We prove that the validity of this procedure (non-asymptotic control of the type I error even under data corruption) and present conditions for consistency (asymptotic control of the type II error). In Sections 3 and 4, we construct robust kernel tests for the two-sample and independence testing frameworks, respectively. For each robust test, we prove minimax optimality under data corruption in terms of the respective MMD and HSIC kernel metrics. In Section 5, we discuss related work on differential privacy which can be leveraged to construct robust tests. We empirically verify the validity of all robust tests in Section 6, and highlight, on corrupted synthetic and real-world data, the significantly higher power achieved by our DC tests compared to alternative private tests of Kim and Schrab, (2023). Section 7 closes the paper with further discussions. We defer the technical proofs of the main results to the appendix.
@@
Notation. Consider two ordered sets $\mathcal{X}_{n}\coloneqq(X_{1},\ldots,X_{n})$ and ${\mathcal{Y}}_{n}\coloneqq(Y_{1},\ldots,Y_{n})$. The Hamming distance $d_{\mathrm{ham}}(\mathcal{X}_{n},{\mathcal{Y}}_{n}):=\sum_{i=1}^{n}\mathds{1}(X_{i}\neq Y_{i})$ is the number of indices for which the entries of the two ordered sets are different. We let $[n]$ denote the set $\{1,\dots,n\}$ and $[n]_{0}\coloneqq\{0,\dots,n\}$. We define $\bm{\Pi}_{n}$ to be the set of all permutations of $[n]$. Given a permutation $\bm{\pi}\in\bm{\Pi}_{n}$, we denote by $\mathcal{X}_{n}^{\bm{\pi}}$ the permuted ordered set $(X_{\bm{\pi}(1)},\ldots,X_{\bm{\pi}(n)})$. For two sequences of real numbers $a_{n}$ and $b_{n}$, we write $a_{n}\lesssim b_{n}$ if $a_{n}\leq Cb_{n}$ for some constant $C>0$, and write $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$. The global sensitivity $\Delta_{T}$ of a statistic $T$ is defined as the maximum difference in the statistic’s output when evaluated on permuted datasets whose entries differ by at most one, that is
@@
$\Delta_{T}\coloneqq\sup_{\bm{\pi}\in\bm{\Pi}_{n}}\sup_{\begin{subarray}{c}\mathcal{X}_{n},{\mathcal{Y}}_{n}\,:\,d_{\mathrm{ham}}(\mathcal{X}_{n},{\mathcal{Y}}_{n})\leq 1\end{subarray}}\bigl{|}T(\mathcal{X}_{n}^{\bm{\pi}})-T({\mathcal{Y}}_{n}^{\bm{\pi}})\bigr{|}.$
@@
Given values $M_{0},\ldots,M_{B}$, the $(1\!\!-\!\!\alpha)$-quantile of the $M_{i}$’s is defined as $q_{1-\alpha}\coloneqq\inf\bigl{\{}t\in\mathbb{R}:\frac{1}{B+1}\sum_{i=0}^{B}\mathds{1}(M_{i}\leq t)\geq 1-\alpha\bigr{\}}$. We let $\mathsf{Laplace}(0,1)$ denote a Laplace distribution with location and scale parameters $(0,1)$. We also let $\mathsf{Gaussian}(\mu,\sigma,d)$ denote a $d$-dimensional distribution with each dimension being independent and following a Gaussian distribution with mean $\mu$ and standard deviation $\sigma$. Similarly, $\mathsf{Geometric}(p,d)$ denotes a $d$-dimensional distribution with each dimension being independent and following a geometric distribution taking values in $\{0,1,2,\ldots\}$ with parameter $p\in(0,1)$.
@@ 2 CONSTRUCTION OF HYPOTHESIS TESTS ROBUST TO DATA CORRUPTION
In Section 2.1, we set the stage and formalise our framework for robust testing in the present of data corruption. In Section 2.2, we introduce a novel procedure (DC) that enables the construction of robust tests under data corruption using any test statistic with finite sensitivity.
@@ 2.1 Robust testing framework under data corruption
Robustness. Hypothesis tests designed for this setting are robust to data corruption: under the null, manipulating up to $r$ samples will not make the test deviate from the null. This can also be thought of as enlarging the null hypothesis as it only needs to hold for at least $n-r$ samples of the data rather than for all $n$ samples. The number $r$ of maximum sample manipulations to be robust to is specified by the user depending on the application. If $r=0$, we recover the standard testing framework. As $r$ increases, the test becomes more robust but less powerful (i.e., there is a trade-off between robustness and power). If $r=n$, the null would never be rejected.
Exchangeability. We restrict our attention to testing frameworks under which the exchangeability assumption holds (Lehmann and Romano,, 2005, Chapter 15.2) in the non-corrupted setting, which means that, under the null, $\widetilde{\mathcal{X}}_{n}$ is exchangeable: for any permutation $\bm{\pi}\in\bm{\Pi}_{n}$, the joint distributions of $\widetilde{\mathcal{X}}_{n}$, and of $\widetilde{\mathcal{X}}_{n}^{\bm{\pi}}$, are the same. For example, the two-sample and independence testing frameworks, serving as primary applications of our proposal, satisfy this assumption.
2.2 DC procedure: Robust test construction against data corruption
We assume the maximum number of corrupted samples $r$ to be fixed apriori. The aim is then to construct a robust test that is valid in the sense that it controls the type I error at the desired level $\alpha$ even when up to $r$ samples are arbitrarily corrupted.
We now propose a procedure to construct tests which are robust to data corruption. First, recall that, given a statistic $T$, the classical permutation test is defined as rejecting the null when $T_{0}>q_{1-\alpha}(T_{0}\ldots,T_{B})$ where $T_{0}=T(\mathcal{X}_{n})$ and $T_{i}=T(\mathcal{X}_{n}^{\bm{\pi}_{i}})$, $i\in[B]$ for $B$ permutations $\bm{\pi}_{1},\ldots,\bm{\pi}_{B}\in\bm{\Pi}_{n}$. Such a test in the usual setting (no data corruption) controls the type I error at level $\alpha$ non-asymptotically (e.g., Hemerik and Goeman,, 2018). For our setting with at most $r$ corrupted samples, we instead define our DC test as rejecting the null when $T_{0}>q_{1-\alpha}(T_{0}\ldots,T_{B})+2r\Delta_{T}$ (see details in Algorithm 1), where $\Delta_{T}$ is the global sensitivity of $T$. We prove that this results in a well-calibrated non-asymptotic test under $r$ data corruption, and provide sufficient conditions to guarantee its consistency.
@@ 2.2 DC procedure: Robust test construction against data corruption
$\lim_{n\rightarrow\infty}\mathbb{P}_{P_{1}}\bigl{(}\,T({\widetilde{\mathcal{X}}_{n}})~{}>~{}T({\widetilde{\mathcal{X}}_{n}}^{\bm{\pi}})+4r\Delta_{T}\bigr{)}=1,$
@@
We highlight that Lemma 1 is applicable to any test statistic with finite global sensitivity $\Delta_{T}$. Moreover, the validity under data corruption holds in any testing framework for which exchangeability holds under the (uncorrupted) null hypothesis. Finally, we stress that knowledge of the maximum number of potential corruptions $r$ is sufficient for all theoretical guarantees in this paper, including Lemma 1, to hold; knowing the exact number of corrupted samples is not required.
We next turn to two-sample and independence kernel testing as specific applications of these general methods, and present more refined results.